### v1

# ConVerSum: A Contrastive Learning based Approach for Data-Scarce Solution of Cross-Lingual Summarization Beyond Direct Equivalents

Abstract
Cross-lingual summarization (CLS) is a sophisticated branch in Natural Language Processing that demands models to accurately translate and summarize articles from different source languages. Despite the improvement of the subsequent studies, This area still needs data-efficient solutions along with effective training methodologies. To the best of our knowledge, there is no feasible solution for CLS when there is no available high-quality CLS data. In this paper, we propose a novel data-efficient approach, ConVerSum, for CLS leveraging the power of contrastive learning, generating versatile candidate summaries in different languages based on the given source document and contrasting these summaries with reference summaries concerning the given documents. After that, we train the model with a contrastive ranking loss. Then, we rigorously evaluate the proposed approach against current methodologies and compare it to powerful Large Language Models (LLMs)- Gemini, GPT 3.5, and GPT 4 proving our model performs better for low-resource languages’ CLS. These findings represent a substantial improvement in the area, opening the door to more efficient and accurate cross-lingual summarizing techniques.
Keywords Cross-lingual  $\cdot$ summarization  $\cdot$ contrastive learning  $\cdot$ candidate summary  $\cdot$ reference summary  $\cdot$ low-resource
1 Introduction
Cross-lingual Summarization (CLS) is the process of producing a summary in a different language from the source language. This can aid in easily acquiring the gist of a foreign language article. This task can be thought of as a fusion of machine translation (MT) and monolingual summarization (MS), two open natural language processing (NLP) challenges that have been researched for decades [1] [2]. CLS is an extremely complex process from the perspective of both data and models. In terms of data, as opposed to MS, the lack of naturally occurring documents in a source language and comparable summaries in several target languages make it difficult to gather large-scale and human-annotated datasets [3] [4]. From the viewpoint of models, CLS necessitates the ability to translate and summarize, making it challenging to provide precise summaries using CLS independently [5]. Although the absence of high-quality large-scale data has recently been addressed [6], the CLS still needs models that take data-efficient solutions into account.
Previous attempts at CLS used pipeline methods such as translate-then-summarize [7] and summarize-then-translate [8]. Due to the necessity for several models, these methods are not only computationally costly but also experience error propagation [9] among models, which diminishes overall performance rendering it inappropriate for the scenario in real life [3]. To tackle this problem, CLS has recently been accomplished by utilizing Sequence-to-Sequence(Seq2Seq) [10] [11] and Transformers-based models [12] to perform the summarization in a single model [9] [13] [6]. The aforementioned achievements have paved the way for the field of CLS research and progressively fueled curiosity in CLS. However, no one proposed any CLS solution when there is no available large-scale CLS corpus to the best of our knowledge. Thus, we wish to investigate the question: Can we develop a CLS model which will be able to generate coherent and informative summaries when no CLS data is available?
In this paper, we propose a model ConVerSum which provides a data-scarce solution for CLS without parallel corpora by leveraging contrastive learning. At first, we generate versatile candidate summaries in different languages using a seq2seq model, use diverse Beam search to explore word sequences, measure the quality of candidate summaries, apply contrastive learning to minimize positive and negative pairs and train the model with a contrastive ranking loss to improve similarity scores. ConVerSum is evaluated using standard metrics and tested on different cross-lingual language pairs and LLMs. Therefore, our main contribution to this paper is as follows:
  • We propose a contrastive learning-based approach for CLS ConVerSum when there is no large-scale parallel corpus of CLS.
  • We design a contrastive ranking loss to train the model to improve the similarity score of candidate summaries concerning document and reference summaries.
  • We carry out a comprehensive comparative study to identify the best-performing approach and prove ConVerSum’s robustness on different cross-lingual pairs and low-resource languages. We evaluate it with state-of-the-art LLMs like GPT 3.5, GPT 4, and Gemini.
2 Related Works
Early CLS work utilized pipeline approaches, dividing CLS into MS and MT sub-tasks and completing them sequentially. These approaches can be classified as summarize- then-translate [14, 8] or translate-then-summarize [7, 15, 16, 17] based on the sequence of completed sub-tasks. End-to-end methods that performed CLS with a single model achieve popularity with the emergence of neural models. These methods can be categorized into four frameworks [18] such as multi-task framewok [9, 5, 19], knowledge-distillation framework [20, 21], resource-enhanced framework [22, 23, 22] and pre-training framework [24, 25].
Contrastive learning has recently successfully emerged in various fields of NLP [26]. The application of contrastive learning for abstractive summarization is a nascent yet promising branch of research [27, 28, 29].
3 ConVerSum
In this section, we elaborate on our proposed approach ConVerSum. At first, candidate summaries are generated in different languages using monolingual languages. We generate candidate summaries inspired by the candidate generation process of SimCLS [27]. From a monolingual document, numerous candidate summaries in several languages are generated using the seq2seq model, mT5 [25]. Because it was extensively trained on a diverse language dataset, it possesses multilingual skills. The model can be easily adjusted to different language contexts. This scenario makes use of the improved many-to-many (m2m) mT5 checkpoint111csebuetnlp/mT5_m2m_crossSum_enhanced. However not all summaries that are generated are of excellent quality, and it’s possible for repeating summaries to be produced. This is addressed using diverse Beam search to generate candidate summaries, exploring various word sequences and producing high-quality summaries.
Then, the generated candidate summaries are evaluated concerning the reference summaries and the given documents.The evaluation function for candidate summaries is initiated with XLM-RoBERTa [30] base model [^2] FacebookAI/xlm-roberta-base for mapping semantically similar text from several languages into a shared embedding space. This procedure captures the meaning and context of the text. Cosine similarity is used to determine the similarity of candidate summaries and reference summaries to the supplied source document. LaSE [6], a scoring method devised exclusively for assessing CLS, is calculated for each candidate summary and ranked in descending order. As a result, the highest-scoring candidate is more similar to the reference summary, while the lowest-scoring candidate is less comparable. Both cosine similarity and LaSE scores are employed to capture many aspects of summary quality, ensuring a full evaluation as well as the correctness and relevancy of the summaries. This is crucial for tasks requiring accurate factual information and contextual knowledge.
Next, we implement contrastive learning objective for our task. The goal of the contrastive learning is to minimize the distance between positive pairs and maximize the distance between negative pairings in order to improve cross-lingual summary generation. Using a contrastive ranking loss function, the model is trained to push negative pairings further away and bring positive pairs closer together. By performing this, it is ensured that the summaries that are produced are relevant to the context and correctly semantically. To achieve this goal, we construct positive and negative pairs. Positive pairings reflect summaries that are more similar to the reference summary concerning documents, implying greater quality. In contrast, negative pairings comprise summaries that are of inferior quality when compared to other candidate summaries and the reference summary. This method aids in discriminating between superior and worse summaries, increasing the model’s ability to generate high-quality summaries.
The model is trained on cross-lingual summary generation with contrastive ranking loss to capture complex connections between the original documents and their related summaries in different languages. It learns to assign higher similarity scores to positive summary pairings and lower scores to negative combinations. The model’s efficacy is assessed using the LaSE and BERTScore. The model’s resilience is tested on the existing CLS dataset and low-resource monolingual datasets, and performance is compared to the most powerful current LLMs.
Figure 1 illustrates the general structure of ConVerSum. Furthermore, we repeat the whole process using Flan-T5 [31] instead of mT5 where we use the checkpoint333google/flan-t5-large to identify which approach performs better. All the details are explained in Section 4.
Figure 1: General Structure of ConVerSum. Here CS, RS, Lang and Doc refer to candidate summary, reference summary, language and document, respectively. For better realization, we assume score1 > score2 > score3 > score4.
4 Experimental Evaluation
4.1 Training Setup
The main computing resource for model training is the NVIDIA GeForce RTX 3090 GPU. We use the broadly recognized deep learning framework PyTorch [^4] https://pytorch.org/ (version 3.7.9) to develop ConVerSum.
4.2 Training Consideration
We have two things in our consideration for training ConVerSum - (i) training utilizing a large-scale monolingual dataset of high-resource languages and (ii) training utilizing a significantly smaller monolingual dataset of low-resource languages to demonstrate the power of contrastive learning in ConVerSum for both large-scale and small-scale datasets. The ConVerSum’s performance is analyzed on a cross-lingual dataset to verify its robustness.
4.3 Datasets
According to the consideration mentioned in Section 4.2. We choose CNN/DailyMail [32, 33] for the high-resource large-scale monolingual dataset and XLSUM [34] for low-resource significantly smaller monolingual datasets. We choose the monolingual Bengali, Thai, Burmese, and Tigrinya language datasets for our task from XLSUM. Finally, we perform a performance analysis of our model on CrossSum [6], a large-scale CLS dataset comprising 1.68 million article-summary samples in 1,500+ language pairs. The dataset for training and verification of the model is shown in Table 1. The dataset statistics for monolingual datasets for training are shown in Table 2.
Table 1: Datasets for Training and Verification
Training Dataset (Monolingual) Verification Dataset (Cross-lingual)
CNN/DailyMail
XLSUM CrossSum
   • Bengali
   • Thai
   • Burmese
   • Tigrinya
Table 2: Training, Validation, and Test Splits for Monolingual Datasets for Training(Rsrc=Resource, Ds=Dataset)
Rsrc-Ds Monolingual Data Train Val Test
high-large CNN/DailyMail 287113 13368 11490
low-small XLSUM-Bengali 8102 1012 1012
low-small XLSUM-Thai 6616 826 826
low-small XLSUM-Burmese 4569 570 570
low-small XLSUM-Tigrinya 5451 681 681
4.4 Baseline Model
We use the mT5 [25] and FLAN-T5 [31] models as baselines for our tasks. These models are selected for their effectively handled multilingual data and demonstrated state- of-the-art performance in tasks related to NLP, such as text summarization.
4.5 Evaluation Metrics
We use LaSE [6] and BERTScore [35]. LaSE is specifically designed for evaluating CLS considering meaning similarity, language confidence and length penalty. BERTScore exploits the pre-trained contextual embeddings from BERT [36] to identify words in candidate and reference sentences based on cosine similarity.
4.6 ConVerSum: Training on High-Resource Large Scale Dataset
This section elaborates on how we develop ConVerSum on CNN/DailyMail Dataset as we discussed in Section 4.2 and 4.3. We utilize mT5 and Flan-T5 separately as they are our baselines for developing ConVerSum. Then, we perform a comparison analysis to know which Transformer assists us in developing the robust model.
To generate candidate summaries in different languages, we use mT5 checkpoint [^5] https://huggingface.co/csebuetnlp/mT5_m2m_crossSum_enhanced available in Huggingface [^6] https://huggingface.co/. For diversity sampling using diverse Beam search, we employ 8 groups, yielding 8 candidate summaries. The maximum length of the generated candidate summaries is specified at 80 words. Using an identical length guarantees uniformity between summaries in different languages. This uniformity makes the evaluation process easier and allows for fair comparisons. Moreover, languages differ in terms of word length and sentence structure. The summary word restriction guarantees all summaries are reasonably equal in terms of their content and depth, independent of language. For validation and test sets, we also generate the candidate summaries. To encode reference summaries, candidate summaries, and documents, we initiate the evaluation function with the XLM-RoBERTa checkpoint [^7] FacebookAI/xlm-roberta-base as it learns the inner representation of 100 languages. Then, we apply the contrastive learning objective described in Section 3. For training, Adam [37] optimizer is used with learning rate scheduling. The batch size is set to 2 for candidate generation and 4 for model training after trying 32, 16, 8 and 4 because of our GPU limitations. This indicates that we always choose the maximum batch size allowed by our GPU memory. We evaluate the performance of ConVerSum on the validation set at every 1000 steps using the LaSE. We set 15 epochs for training.
For experimenting with Flan-T5, we use Flan-T5 checkpoint [^8] https://huggingface.co/google/flan-t5-large to generate candidate summaries in different languages. The other settings like the dataset used for training, and the overall training process with all setups remain as same as described before. This consistent setup makes it easier to do comprehensive research and comparisons across the models of CLS.
4.6.1 Performance Comparison of ConVerSum with Baselines on CrossSum Dataset
To evaluate the performance of ConVerSum and compare it with the baselines mentioned in Section 4.4, we consider three types of combinations - (i) Source English to Target Other Languages, (ii) Source Other Languages to Target English and (iii) Source Other Languages to Target Other Languages Except English and observe the performance of all of these combinations. Table 3, Table 4, Table 5, Table 6, Table 7 and Table 8 demonstrate the performance comparison for these combinations.
Table 3: Performance comparison of (English source and other target) baseline and ConVerSum based on mT5 on CrossSum
Baseline (mT5) ConVerSum
Source-target (English - others) LaSE BERTScore LaSE BERTScore
English - Chinese_simplified 0.5638 0.8573 0.5987 0.8556
English - Hindi 0.6042 0.8510 0.6398 0.8567
English - Bengali 0.5874 0.8407 0.6083 0.8440
English - Japanese 0.5773 0.8364 0.5879 0.8384
English - Marathi 0.6071 0.8433 0.6305 0.8518
English - Thai 0.4121 0.8324 0.4606 0.8323
English - Tigrinya 0.5532 0.8062 0.5740 0.8056
Table 4: Performance comparison of (English source and Others target) baseline and ConVerSum based on Flan-T5 on CrossSum
Baseline (Flan-T5) ConVerSum
Source-target (English to others) LaSE BERTScore LaSE BERTScore
English - chinese_simplified 0.3814 0.8130 0.3890 0.8097
English - Hindi 0.3700 0.8124 0.3707 0.8063
English - Bengali 0.3341 0.7967 0.3328 0.8015
English - Japanese 0.3464 0.7995 0.3570 0.7960
English - Marathi 0.3870 0.8046 0.3750 0.7988
English - Thai 0.2971 0.7944 0.3019 0.7907
English - Tigrinya 0.4352 0.7773 0.3193 0.7722
Table 5: Performance comparison of (Others source and English target) baseline and ConVerSum based on mT5 on CrossSum
Baseline (mT5) ConVerSum
Source-target (others to English) LaSE BERTScore LaSE BERTScore
Chinese_simplified - English 0.4742 0.8418 0.5361 0.8437
Hindi - English 0.5179 0.8335 0.5239 0.8329
Bengali - English 0.5168 0.8263 0.5149 0.8258
Japanese - English 0.5216 0.8341 0.5548 0.8347
Marathi - English 0.5177 0.8290 0.5341 0.8275
Thai - English 0.5078 0.8298 0.5223 0.8315
Tigrinya - English 0.4478 0.7935 0.3758 0.7936
Table 6: Performance comparison of (Others source and English target) baseline and ConVerSum based on Flan-T5 on CrossSum
Baseline (Flan-T5) ConVerSum
Source-target (others to English) LaSE BERTScore LaSE BERTScore
Chinese_simplified - English 0.1181 0.6635 0.1425 0.5891
Hindi - English 0.1332 0.6760 0.1352 0.5537
Bengali - English 0.1292 0.6879 0.1187 0.5593
Japanese - English 0.2168 0.5592 0.2442 0.3927
Marathi - English 0.1278 0.6678 0.1315 0.5815
Thai - English 0.1359 0.6809 0.1748 0.6603
Tigrinya - English 0.1039 0.5217 0.1088 0.3877
Table 7: Performance comparison of (Others source and Others target) baseline and ConVerSum based on mT5 on CrossSum
Baseline (mT5) ConVerSum
Source-target (others to others) LaSE BERTScore LaSE BERTScore
Chinese_traditional -Ukrainian 0.5286 0.8321 0.5624 0.8305
Hindi - Uzbek 0.5217 0.7786 0.4664 0.7804
Punjabi - Telugu 0.5227 0.8419 0.5298 0.8414
Japanese - Burmese 0.5892 0.8493 0.65 0.8613
Marathi - Russian 0.5816 0.8281 0.5769 0.8238
Thai-Tamil 0.5674 0.8584 0.6445 0.8673
Table 8: Performance comparison of (Others source and Others target) baseline and ConVerSum based on Flan-T5 on CrossSum
Baseline (Flan-T5) ConVerSum
Source-target (others to others) LaSE BERTScore LaSE BERTScore
Chinese_traditional - Ukrainian 0.1944 0.6020 0.2106 0.6036
Hindi - Uzbek 0.1588 0.6319 0.1527 0.6920
Punjabi - Telugu 0.2081 0.6389 0.2089 0.5713
Japanese - Burmese 0.2277 0.5264 0.2512 0.4219
Marathi - Russian 0.2060 0.6394 0.1979 0.6310
Thai - Tamil 0.2047 0.5259 0.2245 0.5358
4.6.2 Performance Comparison of ConVerSum and LLM for Low Resource Languages
Evaluating the performance of the state-of-art LLMs on various datasets and numerous NLP tasks is a trending research [38, 39]. Inspired by this trend, in this section, we compare the performance of ConVerSum to cutting-edge LLMs - ChatGPT-3.5, ChatGPT-4o, and Gemini 1.5 Pro to measure our ConVerSum’s efficiency in generating high-quality summaries in a variety of low-resource languages. To assess LLMs’ proficiency, we prompt them to rate their level of confidence between 1 to 10 in generating accurate cross-lingual summaries in 45 different languages. Responses are repeatedly generated to ensure clarity. All LLMs rate their capability as low (1-4) for CLS on low-resource languages. We choose three language pairs mentioned in Table 9 based on their ratings for this experiment.
{mdframed}
[linecolor=black, linewidth=2pt, leftmargin=1em, rightmargin=1em, innertopmargin=1em, innerbottommargin=1em] "How confident are you to generate high-quality cross-lingual summary concisely and informatively for low-resource languages? Find the list of languages - Amharic, Arabic, Azerbaijani, Bengali, Burmese, Chinese_simplified, Chinese_traditional, English, French, Gujarati, Hausa, Hindi, Igbo, Indonesian, Japanese, Kirundi, Korean, Kyrgyz, Marathi, Nepali, Oromo, Pashto, Persian, Pidgin, Portuguese, Punjabi, Russian, Scottish_gaelic, Serbian_cyrillic, Serbian_latin, Sinhala, Somali, Spanish, Swahili, Tamil, Telugu, Thai, Tigrinya, Turkish, Ukrainian, Urdu, Uzbek, Vietnamese, Welsh, Yoruba. Rate your confidence level for cross-lingual summarization on a scale of 1 to 10 for the given languages."
Then, we use manual prompting for LLMs which is free of cost. For our selected language pairs, we provide the following prompt where the source document is provided from CrossSum [6] test set and extract the responses manually from LLMs to measure the LaSE and BERTScore and compare those with ConVerSum which are mentioned in Table 9. {mdframed}[linecolor=black, linewidth=2pt, leftmargin=1em, rightmargin=1em, innertopmargin=1em, innerbottommargin=1em] "Summarize the given text in target language, preferably in 80 words, concisely and informative. .. (source document is given here)"
Table 9: Some Performance Measures on LLM vs Our Proposed Model for Low-resource Languages
Source-Target ChatGPT 3.5 ChatGPT 4o ConVerSum Gemini 1.5 Pro
LaSE BERTScore LaSE BERTScore LaSE BERTScore
Amharic-Igbo
Attempts to
generate
monolingual
summary
0.7264 0.7364 0.6377 0.7836 Unable to summarize
Tigrinya-Arabic 0.2798 0.7481 0.5246 0.7592 0.5241 0.7936
Burmese-Bengali
Attempts to
generate
monolingual
summary.
After a few
prompts,
generate
inaccurate CLS
Input a document
of 699 words and
unable to handle
the long text
0.6030 0.8452
4.7 ConVerSum: Training on Low-Resource Language Datasets
This section emphasizes how we develop ConVerSum for CLS on XLSUM [34] Dataset as we mentioned in Section 4.3 to observe how ConVerSum performs on the low-resource small-scale datasets. We choose Bengali, Thai, Burmese and Tigrinya from the XLSUM dataset. As the results using mT5 outperform Flan-T5 in Section 4.6.1, we choose only using mT5 in the case discussed in this section. Therefore, all training setups are kept the same setup mentioned in Section 4.6.
4.7.1 Performance Comparison of ConVerSum with Baselines on CrossSum Dataset
To evaluate the performance of ConVerSum and compare it with the baselines mentioned in Section 4.4, we consider two directions - source low-resource to target another language and source any language to target low-resource language. Table 10, Table 11, Table 12, and Table 13 demonstrate the performance comparison for these combinations for training ConVerSum on Bengali, Thai, Burmese and Tigrinya, respectively.
Table 10: Performance Comparison of ConVerSum Trained on XLSUM-Bengali with Baseline on CrossSum Dataset
Baseline (mT5) ConVerSum
Source-Target LaSE BERTScore LaSE BERTScore
Bengali-English 0.5168 0.8263 0.5148 0.8257
Bengali-Arabic 0.5650 0.8229 0.5535 0.8205
Bengali-Marathi 0.6197 0.8493 0.6357 0.8545
Ukrainian-Bengali 0.6110 0.8493 0.6114 0.8470
Hausa-Bengali 0.5831 0.8493 0.6846 0.8629
Table 11: Performance Comparison of ConVerSum Trained on XLSUM-Thai with Baseline on CrossSum Dataset
Source-Target Baseline(mT5) ConVerSum
LaSE BERTScore LaSE BERTScore
Thai-English 0.5078 0.8298 0.5223 0.8315
Thai-Bengali 0.5937 0.8445 0.6447 0.8487
Thai-Marathi 0.5998 0.8411 0.6677 0.8465
Uzbek-Thai 0.6672 0.9305 0.6672 0.9390
Amharic-Thai 0.3783 0.8340 0.4310 0.8268
Table 12: Performance Comparison of ConVerSum Trained on XLSUM-Burmese with Baseline on CrossSum Dataset
Source-Target Baseline(mT5) ConVerSum
LaSE BERTScore LaSE BERTScore
Burmese-English 0.5394 0.8287 0.5422 0.8283
Burmese-Bengali 0.6146 0.8400 0.5973 0.8448
Burmese-Marathi 0.4247 0.8371 0.6821 0.8648
Uzbek-Burmese 0.4918 0.8335 0.4794 0.8431
Amharic-Burmese 0.6345 0.8689 0.6054 0.8599
Table 13: Performance Comparison of ConVerSum Trained on XLSUM-Tigrinya with Baselines on CrossSum Dataset
Baseline(mT5) ConVerSum
Source-Target LaSE BERTScore LaSE BERTScore
Tigrinya-English 0.4478 0.7935 0.3758 0.7937
Tigrinya-Arabic 0.4137 0.7837 0.5241 0.7966
Tigrinya-Marathi 0.4185 0.7762 0.6423 0.7916
Chinese_simplified-Tigrinya 0.4390 0.8288 0.5497 0.8288
Somali-Tigrinya 0.2241 0.7829 0.5537 0.7801
5 Result Analysis
5.1 For Training on High-Resource Large-Scale Dataset
From Table 3 we can say that ConVerSum performs better for LaSE for all source-target language pairs we choose which proves the effectiveness of our model in CLS. While calculating BERTScore, three cases - English to Chinese_simplified, Thai and Tigrinya, ConVerSum achieves almost similar performance as the baseline mT5. Table 5 shows that ConVerSum achieves better performance for LaSE for all of the source-target language pairs we experimented with except Bengali to English and Tigrinya to English though these two cases show almost similar performance to the baseline. While calculating BERTScore, three cases - Hindi-English, Bengali-English and Marathi-English, ConVerSum achieves almost identical performance as the baseline mT5. Table 7 illustrates that ConVerSum achieves better performance for LaSE for almost all of the source-target language pairs we experimented with. While calculating BERTScore, for three cases, ConVerSum achieves almost similar performance as the baseline mT5.
From Table 4, Table 6 and Table 8, it is clear that the performance of ConVerSum based on Flan-T5 and baseline Flan-T5 itself performs significantly lower than the ConVerSum of mT5 and also mT5 as a baseline itself for all combinations mentioned in Section 4.6.1.
5.2 For Training on Low-Resource Small-Scale Dataset
From Table 10, Table 11, Table 12 and Table 13 we can wrap up that ConVerSum outperforms baseline for most of the language pairs and only for a few instances, it shows almost identical performance to baseline.
5.3 ConVerSum vs LLMs
Based on our experiments on LLM for low-resource languages demonstrated in Table 9, we have some observations on LLMs. These are as follows:
5.3.1 Gemini 1.5 Pro:
Gemini only supports a few languages and frequently fails to correctly detect the specified source language. It struggles to produce summaries from documents in low-resource source languages and is equally unable to provide summaries in low-resource target languages, demonstrating its shortcomings in multilingual and low-resource language contexts.
5.3.2 ChatGPT 3.5:
When using ChatGPT 3.5 to handle low-resource languages, the model is frequently unable to provide high-quality, accurate summaries. For several low-resource languages, it generates monolingual summaries rather than CLS. Furthermore, the model produces inefficient summaries even after several prompts, showing its difficulties in properly digesting and summarizing content in low-resource language contexts.
5.3.3 ChatGPT 4o:
When reviewing the capabilities of the latest and most powerful LLM, ChatGPT-4o, which was launched on May 13, 2024, it is stated that, while this model is capable of handling low-resource languages, it still has limits. It cannot handle large documents (more than 700 words) for several low-resource languages, such as Burmese, emphasizing areas where further improvements are required.
5.4 LaSE vs BERTScore
Instances, where ConVerSum’s LaSE is higher than the baseline, highlight its strength in cross-lingual summary generation preserving the overall structure with meaning across different languages. Moreover, higher BERTScore focuses on higher semantic similarity. This dual-metric evaluation delivers significant insights to satisfy the demands of different application scenarios.
5.5 Overall Discussion
ConVerSum performs better in both high- and low-resource scenarios, and large- and small-scale datasets, and consistently beats the baseline models. It even performs better than GPT-3.5 and almost as well as GPT-4o in low-resource language settings, outperforming the capabilities of current LLMs. The observation that ConVerSum performs well even with limited training data for low-resource languages highlights the efficacy of the contrastive learning approach used. These outcomes validate the robustness and flexibility of ConVerSum, highlighting the ability of contrastive learning to improve model performance even in resource-constrained scenarios.
6 Conclusion
The paper introduces ConVerSum, a novel contrastive learning-based CLS approach. By extracting and generalizing semantic representations across languages, this technique enhances CLS’s adaptability and durability. A comprehensive comparison with state-of-the-art models and various cross-lingual language pairs proves the effectiveness of the approach. This strategy works especially well when there is no CLS parallel data available, as traditional approaches might not be able to produce accurate findings. ConVerSum improves cross-lingual summaries by utilizing the unique characteristics of contrastive learning, therefore establishing a new standard for further studies. Nonetheless, there are several difficulties, such as the large computational capacity needed for training and the requirement for a large number of candidate summaries from various language sets. In the future, there are potential improvements in performance, adjustments to methods, and the use of guidelines to guarantee fairness and biases.