### v2

# ConVerSum: A Contrastive Learning-based Approach for Data-Scarce Solution of Cross-Lingual Summarization Beyond Direct Equivalents

Abstract.
Cross-lingual summarization (CLS) is a sophisticated branch in Natural Language Processing that demands models to accurately translate and summarize articles from different source languages. Despite the improvement of the subsequent studies, This area still needs data-efficient solutions along with effective training methodologies. To the best of our knowledge, there is no feasible solution for CLS when there is no available high-quality CLS data. In this paper, we propose a novel data-efficient approach, ConVerSum, for CLS leveraging the power of contrastive learning, generating versatile candidate summaries in different languages based on the given source document and contrasting these summaries with reference summaries concerning the given documents. After that, we train the model with a contrastive ranking loss. Then, we rigorously evaluate the proposed approach against current methodologies and compare it to powerful Large Language Models (LLMs)- Gemini, GPT 3.5, and GPT 4o proving our model performs better for low-resource languages’ CLS. These findings represent a substantial improvement in the area, opening the door to more efficient and accurate cross-lingual summarizing techniques.
cross-lingual, summarization, low-resource, LLM, contrastive learning
copyright: noneconference: ; journalyear: ;doi: isbn:
1. Introduction
Cross-lingual Summarization (CLS) is the process of producing a summary in a different target language (e.g. Bengali) from the source language (e.g. Japanese) (Figure 1). This can aid in easily acquiring the gist of a foreign language article. This task can be thought of as a fusion of machine translation (MT) and monolingual summarization (MS), two open natural language processing (NLP) challenges that have been researched for decades (Paice, 1990; Brown et al., 1993). CLS is an extremely complex process from the perspective of both data and models. In terms of data, as opposed to MS, the lack of naturally occurring documents in a source language and comparable summaries in several target languages make it difficult to gather large-scale and human-annotated datasets (Ladhak et al., 2020; Perez-Beltrachini and Lapata, 2021). From the viewpoint of models, CLS necessitates the ability to translate and summarize, making it challenging to provide precise summaries using CLS independently (Cao et al., 2020a). Although the absence of high-quality large-scale data has recently been addressed (Bhattacharjee et al., 2023), the CLS still needs models that take data-efficient solutions into account.
Figure 1. Example of CLS
[…]
Previous attempts at CLS used pipeline methods such as translate-then-summarize (Leuski et al., 2003) and summarize-then-translate (Wan et al., 2010). Due to the necessity for several models, these methods are not only computationally costly but also experience error propagation (Zhu et al., 2019) among models, which diminishes overall performance rendering it inappropriate for the scenario in real life (Ladhak et al., 2020). To tackle this problem, CLS has recently been accomplished by utilizing Sequence-to-Sequence(Seq2Seq) (Cho et al., 2014; Sutskever et al., 2014) and Transformers-based models (Vaswani et al., 2017) to perform the summarization in a single model (Zhu et al., 2019; Cao et al., 2020b; Bhattacharjee et al., 2023). The aforementioned achievements have paved the way for the field of CLS research and progressively fueled curiosity in CLS. However, no one proposed any CLS solution when there is no available large-scale CLS corpus to the best of our knowledge. Thus, we wish to investigate the question: Can we develop a CLS model which will be able to generate coherent and informative summaries when no CLS data is available?
In this paper, we propose a model ConVerSum which provides a data-scarce solution for CLS without parallel corpora by leveraging contrastive learning. At first, we generate versatile candidate summaries in different languages using a seq2seq model, use diverse Beam search to explore word sequences, measure the quality of candidate summaries, apply contrastive learning to minimize positive and maximize negative pairs and train the model with a contrastive ranking loss to improve similarity scores. ConVerSum is evaluated using standard metrics and tested on different cross-lingual language pairs and Large Language Models (LLMs). Therefore, our main contribution to this paper is as follows:
  • We propose a contrastive learning-based approach for CLS ConVerSum when there is no large-scale parallel corpus of CLS.
  • We design a contrastive ranking loss to train the model to improve the similarity score of candidate summaries concerning document and reference summaries.
  • We carry out a comprehensive comparative study to identify the best-performing approach and prove ConVerSum’s robustness on different cross-lingual pairs and low-resource languages. We evaluate it with state-of-the-art LLMs like GPT 3.5, GPT 4o, and Gemini.
2. Related Works
Over the past five years, more than twenty articles have been published by researchers who did their best to solve the CLS task (Wang et al., 2022). Moreover, contrastive learning in abstractive summarization has attracted the attention of researchers lately (Zhang et al., 2022). We analyze existing publications to identify the most sophisticated strategies employed in recent studies. This section provides a full discussion of relevant articles and their working techniques on CLS and contrastive learning in summarization.
2.1. Pipeline Methods
Early CLS work focused on pipeline approaches, which decompose CLS into MS and MT sub-tasks and complete them sequentially. These approaches can be classified as summarize-then-translate (Orăsan and Chiorean, 2008; Wan et al., 2010) or translate-then-summarize (Leuski et al., 2003; Wan, 2011; Boudin et al., 2011; Yao et al., 2015) based on the sequence of completed sub-tasks. The pipeline method is straightforward however has some drawbacks. First of all, they are computationally very expensive for using multiple models; secondly, these techniques are prone to error propagation, resulting in unsatisfactory overall performance (Zhu et al., 2019). Moreover, latency occurs during inference. Additionally, training MT models requires either a big corpus or commercial services.
2.2. End-to-End Methods
Advancements in seq2seq (Cho et al., 2014; Sutskever et al., 2014) and Transformer-based models (Vaswani et al., 2017) have led to the invention of end-to-end methods for performing CLS with a single model to resolve the drawbacks mentioned in Section 2.1. The end-to-end methods can be classified into four categories - multi-task framework, knowledge distillation framework, resource-enhanced framework and pre-training framework.
Multi-task is a machine learning approach in which a single model is designed by combining several related tasks. Several studies train unified models using CLS in conjunction with related tasks such as MT and MS (Zhu et al., 2019; Cao et al., 2020a; Takase and Okazaki, 2020; Bai et al., 2021; Liang et al., 2022). CLS models thus gain from the related tasks as well.
Knowledge distillation is a machine learning approach used to reduce a large, complicated model to a smaller, simpler one. This is accomplished by transferring information from the larger model, known as the teacher model, to the smaller model, known as the student model (Hinton et al., 2015). To teach the CLS model in the knowledge-distillation framework, several researchers employ either MT/MS or both models because to have a mutual connection. In addition to CLS labels, the student model can gain knowledge from the output or hidden state of the teacher’s models. Extensive MS and MT datasets are leveraged to train MS and MT models, respectively which are bidirectional GRU (Cho et al., 2014). These models act as teachers for a CLS student model, with KL-divergence aligning generation probabilities and maximizing the training objective (Ayana et al., 2018). Afterward, transformer (Vaswani et al., 2017) is used as the backbone for the CLS student model and the MS teacher model and further trains the student model (Duan et al., 2019a).
The resource-enhanced framework enriches input documents with extra resources. Output summaries are generated based on encoded and enriched information. Translation patterns in CLS are investigated by encoding source language texts using a transformer encoder (Zhu et al., 2020). The TextRank toolset (Mihalcea and Tarau, 2004) helps create article graphs by extracting important information from input sequences (Jiang et al., 2022). A transformer-based clue encoder and a graph neural network-based graph encoder are used, respectively, to encode these clues and article graphs. Ultimately, the translation distribution from (Zhu et al., 2020) is included in final summaries generated by a transformer decoder with dual cross-attention methods.
In the context of NLP, pre-trained models are neural network architectures that are extensively trained on textual data to identify linguistic patterns and semantic representations. These models are trained on extensive text corpora, which helps them understand word and phrase semantic meanings as well as contextual relationships. By using pre-trained NLP models, transfer learning can be accomplished more successfully. This is because the gained information can be fine-tuned on certain NLP tasks improving efficiency and performance. The pre-trained models revolutionize NLP (Qiu et al., 2020).
Recently, multi-lingual pre-trained generative models have performed exceptionally well in various multi-lingual NLP tasks. mBART (Liu et al., 2020) is a multilingual pre-trained model which is another variant of BART (Lewis et al., 2020) model developed by Facebook AI. mBART is pre-trained with denoising objectives based on the BART architecture and vast unlabeled multilingual datasets. Specifically, it is trained in over 25 languages text data. Initially excelling in MT (Liu et al., 2020), a recent study shows that it can outperform many multi-task CLS models on large-scale datasets with straightforward fine-tuning (Liang et al., 2022). Subsequently, the language processing capabilities of mBART are expanded from 25 to 50 languages with mBART-50 (Tang et al., 2021). Aside from the BART-style pre-trained models, mT5 (Xue et al., 2021) shines out as a multilingual version of T5 (Raffel et al., 2020), trained across 101 languages with the T5-style span corruption objective. Despite the remarkable performance of these general pre-trained models, they lack explicit cross-lingual supervision and instead rely only on denoising or span corruption objectives across several languages. They have thus far not fully explored their cross-lingual potential.
2.3. Contrastive Learning for Monolingual Abstractive Summarization
Contrastive learning has recently successfully emerged in various fields of NLP (Zhang et al., 2022), to learn good and discriminative representations. Contrastive learning is a self-supervised approach to training models to differentiate between samples that are similar versus dissimilar. This approach is popularized in the SimCLR framework (Chen et al., 2020), which uses contrastive loss to minimize the distance between representations for similar data points and maximize it for dissimilar points.
The application of contrastive learning for abstractive summarization is a nascent yet promising branch of research. The SimCLS framework is designed to bridge the gap between standard seq2seq models and learning objectives by utilizing contrastive learning to adapt to evaluation metrics with the learning objective (Liu and Liu, 2021). A contrastive attention mechanism is proposed for abstractive sentence summarization that builds on the seq2seq framework by combining conventional attention, which focuses on relevant parts of the source sentence, and opponent attention, which focuses on less relevant parts, with both trained in opposite directions using novel softmax and softmin functions (Duan et al., 2019b). COLO, a contrastive learning-based re-ranking framework for one-stage summarizing is introduced that aligns training and assessment by producing summaries based on summary-level scores without the requirement for extra modules or parameters (An et al., 2022). CALMS, a unified summarizing model is constructed which is trained with contrastive learning by using MLGSum, a large-scale multilingual summarizing corpus including 1.1 million articles and summaries in 12 languages (Wang et al., 2021). It incorporates contrastive sentence ranking and sentence-aligned replacement to improve extractive skills and align sentence-level representations across languages.
3. ConVerSum
In this section, we elaborate on our proposed approach ConVerSum which is categorized into several phases. Figure 2 illustrates the general structure of ConVerSum. ConVerSum deals with the data scarcity challenges where the existing models require high-quality large-scale datasets as discussed in Section 2.
[…]
Figure 2. General Structure of ConVerSum. Here CS, RS, Lang and Doc refer to candidate summary, reference summary, language and document, respectively. For better realization, we assume score1 ¿ score2 ¿ score3 ¿ score4 …scoreN.
3.1. First Phase: Candidate Summary Generation in Different Languages
In this phase, candidate summaries are generated in different languages using monolingual languages. We generate candidate summaries inspired by the candidate generation process of SimCLS (Liu and Liu, 2021). Though SimCLS is developed for monolingual abstractive summarization, their candidate generation process motivates us to create a robust CLS model. The detailed procedure is as follows.
3.1.1. Seq2Seq Model Selection
A renowned seq2seq model, mT5 (Xue et al., 2021) is utilized to generate multiple candidate summaries in different languages from the given monolingual document. mT5 has multilingual capabilities as it is pre-trained on a vast and diversified multilingual dataset spanning more than 100 languages. This thorough training enables the model to create summaries in many languages from a single monolingual source, making it very versatile and adaptable to a variety of linguistic settings. The enhanced many-to-many(m2m) mT5 checkpoint111csebuetnlp/mT5_m2m_crossSum_enhanced fine-tuned on all cross-lingual pairs of CrossSum (Bhattacharjee et al., 2023) dataset is employed in this case.
3.1.2. Diverse Summary Exploration
Many candidate summaries can be generated by utilizing the seq2seq model as described in Section 3.1.1. However, not all generated candidate summaries are high-quality. Moreover, there is a possibility of generating repetitive summaries and all summaries in the same language. To tackle these issues, diverse Beam search (Vijayakumar et al., 2016) is used during the generation of candidate summaries which explores different sequences of words in the generated summary, contributes to generating diverse candidate summaries, and finds out the high-quality summaries.
3.2. Second Phase: Candidate Summary Quality Measurement
In this phase, the generated candidate summaries as described in Section 3.1 are measured concerning the reference summaries and the given documents. The detailed process is as follows:
3.2.1. Evaluation Function Initialization
The evaluation function is initiated using XLM-RoBERTa (Conneau et al., 2020). The base model of XLM-RoBERTa [^2] FacebookAI/xlm-roberta-base available in huggingface transformers library [^3] https://huggingface.co/ is specifically employed in this case. The purpose of utilizing XLM-RoBERTa is to map semantically similar text from various languages into a shared embedding space. Encoding the candidate summaries, reference summaries, and documents can capture the meaning and context of the text.
3.2.2. Similarity Calculation
For similarity calculation, cosine similarity is utilized (Rahutomo et al., 2012). The purpose of calculating similarity is to identify how similar the candidate summaries and the reference summary are to the given source document. The general representation of the similarities between the source document, reference summary, and candidate summaries is as follows where C, R and S represent the candidate summaries, reference summary, and the source document, respectively.
$\text{Similarity}=f(C,R,S)=\frac{C\cdot S+R\cdot S}{\|C\|\|S\|+\|R\|\|S\|}$
This score is utilized to measure the contrastive ranking loss described in Section 3.3.2.
3.2.3. Candidate Summary Scoring
While generating candidate summaries as described in Section 3.1, LaSE (Bhattacharjee et al., 2023), which is specifically designed for evaluating CLS is computed concerning the reference summary. That means a LaSE score is assigned to each candidate summary and sorted in descending order. This indicates the highest-scoring candidate is more similar to the reference summary and the lowest-scoring candidate is less similar to the reference summary. This process can be defined as follows where $C_{i}$ and R represent the i-th candidate summary and reference summary, respectively.
$\text{LaSE}(S_{1})\geq\text{LaSE}(S_{2})\geq\dots\geq\text{LaSE}(S_{n})=f(C_{i},R)\quad\text{for each}\ i\in\{1,2,\dots,n\}$
Both cosine similarity and LaSE scores are used in the evaluation of candidate summaries to capture various characteristics of summary quality. Every metric offers distinct perspectives that enhance one another, resulting in a more thorough assessment as there is a possibility of generating bias using only one scoring method as CLS is an extremely complex task. Striking a balance between semantic coherence and exact content overlap enables a thorough examination and guarantees that the summaries that are produced are correct and relevant. This thorough method produces summaries that are of higher quality and more dependability, which is important for tasks requiring accurate factual information in addition to contextual knowledge.
3.3. Third Phase: Contrastive Learning Approach
We apply a contrastive learning objective that cultivates a common representation space to improve the quality of cross-lingual summary generation. The objective of this approach is to maximize the distance between negative pairings and minimize the distance between positive pairs. The model is trained to push negative pairings farther apart and bring positive pairs closer together in the representation space by employing a contrastive ranking loss function. This ensures that the generated summaries are both semantically correct and pertinent to the given context. The model’s capacity to differentiate between and provide excellent summaries is much enhanced by this approach. In this section, the whole process is described elaborately.
3.3.1. Positive-Negative Pair Construction
To implement the contrastive learning objective, we need to construct the pairs that are used for positive and negative samples for contrastive learning. To properly evaluate the quality of candidate summaries, we use a positive-negative pair creation procedure. This entails computing the scores of candidate summaries for the provided documents. Positive pairings include candidate summaries that are more comparable to the reference summary, implying greater quality. In contrast, negative pairings comprise summaries that are of inferior quality when compared to other candidate summaries and the reference summary. This strategy aids in discriminating between better and inferior summaries, hence improving the model’s capacity to produce high-quality summaries.
3.3.2. Contrastive Ranking Loss
The goal of the Contrastive Ranking Loss approach is to maximize the correlations between rankings of various candidate summaries. This is accomplished by adding a margin, which acts as a threshold to control the amount of space between summaries in pairs. This margin makes sure that summaries that should be similar are represented by positive pairings, which are encouraged to have high similarity scores, while summaries that should be dissimilar are represented by negative pairs, which are encouraged to have low similarity scores. The approach efficiently separates pertinent and extraneous summaries by enforcing this margin. Furthermore, Contrastive Ranking Loss promotes proper ranking connections by punishing deviations from the intended ranking order. This means that summaries should be ordered based on their relevance or similarity to an expected reference point. By optimizing the loss function using these ranking connections, the technique guarantees that the most relevant summaries are prioritized in the output. This loss can be mathematically represented as follows:
$\mathcal{L}=\sum_{i=1}^{n}\max(0,\text{margin}-f(C_{i},R^{+})+f(C_{i},R^{-}))$
where, $R^{+}$ and $R^{-}$ represent the positive and negative pairs, respectively.
3.4. Fourth Phase: Model Training
The model is trained on the cross-lingual summary generation task by using the contrastive ranking loss described in Section 3.3.2. This process entails modifying the model’s parameters to successfully capture the complex connections between the source documents and their related summaries across multiple languages. This is accomplished by using Contrastive Loss as the optimization goal, which provides a margin to encourage the model to differentiate between important and irrelevant summary pairs. During training, the model learns to give high similarity scores to positive summary pairings that transmit comparable information across languages and low similarity scores to negative pairs that include different documents.
3.5. Fifth Phase: Fine-Tuning and Evaluation
By fine-tuning the model, it obtains the capacity to create short and informative summaries that keep the substance of the original material across language limitations, thereby increasing the performance of cross-lingual summarizing tasks. Then, the model’s effectiveness is evaluated by the standard evaluation metrics LaSE (Bhattacharjee et al., 2023) and BERTScore (Zhang* et al., 2020).
3.6. Sixth Phase: Model Verification
As we developed our model on a monolingual high-resource dataset, we need to verify the model’s robustness on the existing CLS dataset. Furthermore, we explore our approach to different low-resource monolingual datasets to ensure adaptability in scenarios when there is no parallel CLS data. Moreover, we make a performance comparison of our model to the most powerful latest LLMs.
Furthermore, we repeat the whole process using Flan-T5 (Chung et al., 2022) instead of mT5 where we use the checkpoint444google/flan-t5-large to identify which approach performs better. All the details are explained in Section 4.
4. Experimental Evaluation
4.1. Training Setup
The main computing resource for model training is the NVIDIA GeForce RTX 3090 GPU. We use the broadly recognized deep learning framework PyTorch [^5] https://pytorch.org/ (version 3.7.9) to develop ConVerSum.
4.2. Training Consideration
We have two things in our consideration for training ConVerSum. The first is to train ConVerSum utilizing a large-scale monolingual dataset of high-resource languages. The second is to train ConVerSum utilizing a significantly smaller monolingual dataset of low-resource languages. The reason behind considering these two options is to demonstrate how ConVerSum works while training the model on a large-scale dataset of high-resource language and a significantly smaller dataset of low-resource language which proves the power of contrastive learning for both large-scale and small-scale datasets. For both cases, we analyze the performance of ConVerSum on a cross-lingual dataset to verify the model’s robustness.
4.3. Datasets
According to the consideration mentioned in Section 4.2. We choose CNN/DailyMail (Hermann et al., 2015; Nallapati et al., 2016) for the high-resource large-scale monolingual dataset and XLSUM (Hasan et al., 2021) for low-resource significantly smaller monolingual datasets. We choose the monolingual Bengali, Thai, Burmese, and Tigrinya language datasets for our task from XLSUM. Finally, we perform a performance analysis of our model on CrossSum (Bhattacharjee et al., 2023), a large-scale CLS dataset comprising 1.68 million article-summary samples in 1,500+ language pairs. The dataset for training and verification of the model is shown in Table 1. The dataset statistics for monolingual datasets for training are shown in Table 2. The dataset splits are publicly available and provided by the author and intended for efficient training, validation, and unbiased testing to ensure robust model performance across various languages and summarization tasks.
Table 1. Datasets for Training and Verification
Training Dataset (Monolingual) Verification Dataset (Cross-lingual)
CNN/DailyMail
XLSUM CrossSum
   • Bengali
   • Thai
   • Burmese
   • Tigrinya
Table 2. Training, Validation, and Test Splits for Monolingual Datasets for Training
Resource-Dataset Monolingual Data Train Val Test
high-large CNN/DailyMail 287113 13368 11490
low-small XLSUM-Bengali 8102 1012 1012
low-small XLSUM-Thai 6616 826 826
low-small XLSUM-Burmese 4569 570 570
low-small XLSUM-Tigrinya 5451 681 681
4.4. Baseline Model
We use the mT5 (Xue et al., 2021) and FLAN-T5 (Chung et al., 2022) models as baselines for our tasks. These models are selected for their effectively handled multilingual data and demonstrated state- of-the-art performance in tasks related to NLP, such as text summarization.
4.5. Evaluation Metrics
We use LaSE (Bhattacharjee et al., 2023) and BERTScore (Zhang* et al., 2020). LaSE is specifically designed for evaluating CLS considering meaning similarity, language confidence and length penalty. BERTScore exploits the pre-trained contextual embeddings from BERT (Devlin et al., 2019) to identify words in candidate and reference sentences based on cosine similarity.
4.6. First Consideration: Training ConVerSum on High-Resource Large Scale Dataset
This section elaborates on how we develop ConVerSum on CNN/DailyMail dataset as we discussed in Section 4.2 and 4.3. We utilize mT5 and Flan-T5 separately as they are our baselines for developing ConVerSum. Then, we perform a comparison analysis to know which Transformer assists us in developing the robust model.
To generate candidate summaries in different languages, we use mT5 checkpoint [^6] https://huggingface.co/csebuetnlp/mT5_m2m_crossSum_enhanced available in Huggingface [^7] https://huggingface.co/. For diversity sampling using diverse Beam search, we employ 8 groups, yielding 8 candidate summaries. The maximum length of the generated candidate summaries is specified at 80 words. Using an identical length guarantees uniformity between summaries in different languages. This uniformity makes the evaluation process easier and allows for fair comparisons. Moreover, languages differ in terms of word length and sentence structure. The summary word restriction guarantees all summaries are reasonably equal in terms of their content and depth, independent of language. For validation and test sets, we also generate the candidate summaries. To encode reference summaries, candidate summaries, and documents, we initiate the evaluation function with the XLM-RoBERTa checkpoint [^8] FacebookAI/xlm-roberta-base as it learns the inner representation of 100 languages. Then, we apply the contrastive learning objective described in Section 3. For training, Adam (Bengio and LeCun, 2015) optimizer is used with learning rate scheduling. The batch size is set to 2 for candidate generation and 4 for model training after trying 32, 16, 8 and 4 because of our GPU limitations. This indicates that we always choose the maximum batch size allowed by our GPU memory. We evaluate the performance of ConVerSum on the validation set at every 1000 steps using the LaSE. We set 15 epochs for training.
For experimenting with Flan-T5, we use Flan-T5 checkpoint [^9] https://huggingface.co/google/flan-t5-large to generate candidate summaries in different languages. The other settings like the dataset used for training, and the overall training process with all setups remain as same as described before. This consistent setup makes it easier to do comprehensive research and comparisons across the models of CLS.
4.7. Experimental Results on High-Resource Large Scale Dataset
This section conveys the experimental findings from the analysis of the ConVerSums along with baselines for the high-resource large-scale dataset. The models’ performance is evaluated using standard metrics like LaSE and BERTScore, which serve as a complete assessment of the quality of the generated summaries. There is no significant difference between the training time for the baseline and ConVerSum. The following sections describe the experimental results, providing insights into the proposed methodologies’ effectiveness and reliability for high-resource large-scale dataset.
4.7.1. Model Performance on CNN/DailyMail Test Set
The performance of ConVerSum based on both mT5 and Flan-T5 on the test set of the CNN/DailyMail dataset is shown in Table 3 and Table 4, respectively.
Table 3. Performance of baseline and ConVerSum based on mT5 on the Test set. When we generate candidate summaries using mT5, we choose mT5 as the baseline model to compare our model with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Evaluation
Score
Baseline (mT5)
ConVerSum
on mT5
LaSE 0.3886 0.4408
BERTScore 0.8315 0.8363
Table 4. Performance of baseline and ConVerSum based on Flan-T5 on the Test set. When we generate candidate summaries using Flan-T5, we choose Flan-T5 as the baseline model to compare our model with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Evaluation
Score
Baseline (Flan T5)
ConVerSum
on Flan T5
LaSE 0.3585 0.4012
BERTScore 0.8411 0.8381
In comparison to the mT5 and Flan-T5 baselines, ConVerSum performs better, as demonstrated by the findings shown in Table 3 and Table 4. The ConVerSum’s BERTScore on Flan-T5 is almost identical to the baseline’s, however, the overall performance metrics show a distinct benefit. Furthermore, ConVerSum regularly performs better than the Flan-T5 when compared to the mT5.
4.7.2. Performance Comparison of ConVerSum with Baselines on CrossSum Dataset
To evaluate the performance of ConVerSum and compare it with the baselines mentioned in Section 4.4, we consider three types of combinations - (i) Source English to Target Other Languages, (ii) Source Other Languages to Target English and (iii) Source Other Languages to Target Other Languages Except English and observe the performance of all of these combinations. Table 5, Table 6, Table 7, Table 8, Table 9 and Table 10 demonstrate the performance comparison for these combinations.
Table 5. Performance comparison of (English source and other target) baseline and ConVerSum based on mT5 on CrossSum. When we generate candidate summaries using mT5, we choose mT5 as the baseline model to compare our model with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Baseline (mT5) ConVerSum
Source-target (English - others) LaSE BERTScore LaSE BERTScore
English - Chinese_simplified 0.5638 0.8573 0.5987 0.8556
English - Hindi 0.6042 0.8510 0.6398 0.8567
English - Bengali 0.5874 0.8407 0.6083 0.8440
English - Japanese 0.5773 0.8364 0.5879 0.8384
English - Marathi 0.6071 0.8433 0.6305 0.8518
English - Thai 0.4121 0.8324 0.4606 0.8323
English - Tigrinya 0.5532 0.8062 0.5740 0.8056
Key Findings: ConVerSum outperforms the baseline for LaSE, ensuring CLS robustness. Besides, ConVerSum performs better/almost similar to the baseline for BERTScore emphasizing the meaning similarity.
Table 6. Performance comparison of (English source and Others target) baseline and ConVerSum based on Flan-T5 on CrossSum. When we generate candidate summaries using Flan-T5, we choose Flan-T5 as the baseline model to compare our model with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Baseline (Flan-T5) ConVerSum
Source-target (English to others) LaSE BERTScore LaSE BERTScore
English - Chinese_simplified 0.3814 0.8130 0.3890 0.8097
English - Hindi 0.3700 0.8124 0.3707 0.8063
English - Bengali 0.3341 0.7967 0.3328 0.8015
English - Japanese 0.3464 0.7995 0.3570 0.7960
English - Marathi 0.3870 0.8046 0.3750 0.7988
English - Thai 0.2971 0.7944 0.3019 0.7907
English - Tigrinya 0.4352 0.7773 0.3193 0.7722
Key Findings: ConVerSum performs better/almost similar to the baseline for LaSE. Besides, ConVerSum performs almost similarly to the baseline for BERTScore. However, it has significantly lower performance than ConVerSum on mT5(mentioned in Table 5) as the generated candidates fail to show diverse quality and languages for Flan-T5.
Table 7. Performance comparison of (Others source and English target) baseline and ConVerSum based on mT5 on CrossSum. When we generate candidate summaries using mT5, we choose mT5 as the baseline model to compare our model with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Baseline (mT5) ConVerSum
Source-target (others to English) LaSE BERTScore LaSE BERTScore
Chinese_simplified - English 0.4742 0.8418 0.5361 0.8437
Hindi - English 0.5179 0.8335 0.5239 0.8329
Bengali - English 0.5168 0.8263 0.5149 0.8258
Japanese - English 0.5216 0.8341 0.5548 0.8347
Marathi - English 0.5177 0.8290 0.5341 0.8275
Thai - English 0.5078 0.8298 0.5223 0.8315
Tigrinya - English 0.4478 0.7935 0.3758 0.7936
Key Findings: ConVerSum shows significantly improved performance for Chinese_simplified-English. However, it shows lower than baseline performance for Tigrinya-English. Moreover, our model shows almost similar to baseline performance for Bengali-English as the generated candidates fail to show diverse quality summaries in Tigrinya and Bengali. Besides, ConVerSum outperforms the baseline for the rest of the pairs.
Table 8. Performance comparison of (Others source and English target) baseline and ConVerSum based on Flan-T5 on CrossSum. When we generate candidate summaries using Flan-T5, we choose Flan-T5 as the baseline model to compare our model with it. LaSE, designed specifically for CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Baseline (Flan-T5) ConVerSum
Source-target (others to English) LaSE BERTScore LaSE BERTScore
Chinese_simplified - English 0.1181 0.6635 0.1425 0.5891
Hindi - English 0.1332 0.6760 0.1352 0.5537
Bengali - English 0.1292 0.6879 0.1187 0.5593
Japanese - English 0.2168 0.5592 0.2442 0.3927
Marathi - English 0.1278 0.6678 0.1315 0.5815
Thai - English 0.1359 0.6809 0.1748 0.6603
Tigrinya - English 0.1039 0.5217 0.1088 0.3877
Key Findings: ConVerSum outperforms baseline for LaSE except Bengali-English However, it shows significantly lower performance than ConVerSum on mT5 (mentioned in Table 7) as the generated candidates fail to show diverse quality and languages for Flan-T5.
Table 9. Performance comparison of (Others source and Others target) baseline and ConVerSum based on mT5 on CrossSum. When we generate candidate summaries using mT5, we choose mT5 as the baseline model to compare our model with it. LaSE, designed specifically for CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Baseline (mT5) ConVerSum
Source-target (others to others) LaSE BERTScore LaSE BERTScore
Chinese_traditional -Ukrainian 0.5286 0.8321 0.5624 0.8305
Hindi - Uzbek 0.5217 0.7786 0.4664 0.7804
Punjabi - Telugu 0.5227 0.8419 0.5298 0.8414
Japanese - Burmese 0.5892 0.8493 0.6500 0.8613
Marathi - Russian 0.5816 0.8281 0.5769 0.8238
Thai-Tamil 0.5674 0.8584 0.6445 0.8673
Key Findings: ConVerSum shows significantly improved performance for Japanese-Burmese, and Thai-Tamil than the baseline. However, it shows lower than baseline performance for Hindi-Uzbek, and Marathi-Russian as the generated candidates fail to show diverse quality summaries in Hindi-Uzbek, and Marathi-Russian. Otherwise, ConVerSum outperforms the baseline for the rest of the pairs.
Table 10. Performance comparison of (Others source and Others target) baseline and ConVerSum based on Flan-T5 on CrossSum. When we generate candidate summaries using Flan-T5, we choose Flan-T5 as the baseline model to compare our model with it. LaSE, designed specifically for CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Baseline (Flan-T5) ConVerSum
Source-target (others to others) LaSE BERTScore LaSE BERTScore
Chinese_traditional - Ukrainian 0.1944 0.6020 0.2106 0.6036
Hindi - Uzbek 0.1588 0.6319 0.1527 0.6920
Punjabi - Telugu 0.2081 0.6389 0.2089 0.5713
Japanese - Burmese 0.2277 0.5264 0.2512 0.4219
Marathi - Russian 0.2060 0.6394 0.1979 0.6310
Thai - Tamil 0.2047 0.5259 0.2245 0.5358
Key Findings: ConVerSum outperforms the baseline for LaSE except for Hindi-Uzbek, and Marathi-Russian. However, it shows significantly lower performance than ConVerSum on mT5 (mentioned in Table 9) as the generated candidates fail to show diverse quality and languages for Flan-T5.
4.8. Second Consideration: Training ConVerSum on Low-Resource Language Small Datasets
This section emphasizes how we develop ConVerSum for CLS on XLSUM (Hasan et al., 2021) Dataset as we mentioned in Section 4.3 to observe how ConVerSum performs on the low-resource small-scale datasets. We choose Bengali, Thai, Burmese and Tigrinya from the XLSUM dataset. As the results using mT5 outperform Flan-T5 in Section 4.7.2, we choose only using mT5 in the case discussed in this section. Therefore, all training setups are kept the same setup mentioned in Section 4.6.
4.9. Experimental Results on Low-Resource Small Scale Dataset
4.9.1. Model Performance on XLSUM Test Set
The performance of ConVerSum based on mT5 for training on four separate low-resource languages - Bengali, Burmese, Thai and Tigrinya on the test set of the XLSUM dataset is shown in Table 11.
Table 11. Performance of baseline and ConVerSum based on mT5 on the Test set of XLSUM for four separate languages-Bengali, Burmese, Thai, Tigrinya. As we generate candidate summaries using mT5, we choose mT5 as the baseline model to compare ConVerSum with it. LaSE, designed for specifically CLS evaluation, illustrates the model’s robustness on CLS, which is the best choice for CLS evaluation, and BERTScore illustrates the semantic similarity used as a supplementary metric.
Training Datasets
Evaluation
Metrics
Baseline(mT5) ConVerSum Findings
LaSE 0.5872 0.6228
XLSUM-Bengali BERTScore 0.8658 0.8690
LaSE 0.5579 0.5805
XLSUM-Burmese BERTScore 0.8759 0.8798
LaSE 0.4384 0.5055
XLSUM-Thai BERTScore 0.8475 0.8472
LaSE 0.5265 0.5436
XLSUM-Tigrinya BERTScore 0.8747 0.8768 ConVerSum outperforms baselines for all cases
4.9.2. Performance Comparison of ConVerSum Trained on XLSUM-Bengali with Baseline on CrossSum Dataset
To evaluate the performance of ConVerSum and compare it with the baselines mentioned in Section 4.4, we consider two directions - source low-resource to target another language and source any language to target low-resource language. Table 12, Table 13, Table 14, and Table 15 demonstrate the performance comparison for these combinations for training ConVerSum on Bengali, Thai, Burmese and Tigrinya, respectively.
Table 12. Performance Comparison of ConVerSum Trained on XLSUM-Bengali with Baseline on CrossSum Dataset
Baseline (mT5) Proposed Model
Source-Target LaSE BERTScore LaSE BERTScore
Bengali-English 0.5168 0.8263 0.5148 0.8257
Bengali-Arabic 0.5650 0.8229 0.5535 0.8205
Bengali-Marathi 0.6197 0.8493 0.6357 0.8545
Ukrainian-Bengali 0.6110 0.8493 0.6114 0.8470
Hausa-Bengali 0.5831 0.8493 0.6846 0.8629
Key Findings: ConVerSum shows significant improvement in performance for Hausa-Bengali though lower than baseline performance for Bengali-Arabic. However, it shows almost similar to the baseline performance for Bengali-English. The generated candidates fail to show diverse quality summaries for Arabic. ConVerSum outperforms baseline for the rest of the pairs.
Table 13. Performance Comparison of ConVerSum Trained on XLSUM-Thai with Baseline on CrossSum Dataset
Source-Target Baseline(mT5) Proposed Model
LaSE BERTScore LaSE BERTScore
Thai-English 0.5078 0.8298 0.5223 0.8315
Thai-Bengali 0.5937 0.8445 0.6447 0.8487
Thai-Marathi 0.5998 0.8411 0.6677 0.8465
Uzbek-Thai 0.6672 0.9305 0.6672 0.9390
Amharic-Thai 0.3783 0.8340 0.4310 0.8268
Key Findings: ConVerSum outperforms baseline for all of the pairs and shows significant improvement for Thai-Bengali, Thai-Marathi, Amharic-Thai.
Table 14. Performance Comparison of ConVerSum Trained on XLSUM-Burmese with Baseline on CrossSum Dataset
Source-Target Baseline(mT5) Proposed Model
LaSE BERTScore LaSE BERTScore
Burmese-English 0.5394 0.8287 0.5422 0.8283
Burmese-Bengali 0.6146 0.8400 0.5973 0.8448
Burmese-Marathi 0.4247 0.8371 0.6821 0.8648
Uzbek-Burmese 0.4918 0.8335 0.4794 0.8431
Amharic-Burmese 0.6345 0.8689 0.6054 0.8599
Key Findings: ConVerSum shows significant improvement than baseline performance for Burmese-Marathi, both are low-resource languages and also outperform the Burmese-English. Moreover, our model shows better semantically similar performance than the baseline for Burmese-Bengali and Uzbek-Burmese. However, our model performs lower than the baseline only for Amharic-Burmese.
Table 15. Performance Comparison of ConVerSum Trained on XLSUM-Tigrinya with Baselines on CrossSum Dataset
Baseline(mT5) Proposed Model
Source-Target LaSE BERTScore LaSE BERTScore
Tigrinya-English 0.4478 0.7935 0.3758 0.7937
Tigrinya-Arabic 0.4137 0.7837 0.5241 0.7966
Tigrinya-Marathi 0.4185 0.7762 0.6423 0.7916
Chinese_simplified-Tigrinya 0.4390 0.8288 0.5497 0.8288
Somali-Tigrinya 0.2241 0.7829 0.5537 0.7801
Key Findings: ConVerSum shows significant improvement over baseline performance for all pairs except Tigrinya-English.
4.10. Performance Comparison of ConVerSum and LLMs for Low Resource Languages by Prompting
Evaluating the performance of the state-of-art LLMs on various datasets and numerous NLP tasks is a trending research (Laskar et al., 2023; Kabir et al., 2023). Inspired by this trend, in this section, we compare the performance of ConVerSum to cutting-edge LLMs - ChatGPT-3.5, ChatGPT-4o, and Gemini 1.5 Pro to measure ConVerSum’s efficiency in generating high-quality summaries in a variety of low-resource languages. To assess LLMs’ proficiency, we prompt them to rate their level of confidence between 1 to 10 in generating accurate cross-lingual summaries in 45 different languages. Responses are repeatedly generated to ensure clarity. All LLMs rate their capability as low (1-4) for CLS on low-resource languages. We choose three language pairs mentioned in Table 16 based on their ratings for this experiment.
”How confident are you to generate high-quality cross-lingual summary concisely and informatively for low-resource languages? Find the list of languages - Amharic, Arabic, Azerbaijani, Bengali, Burmese, Chinese_simplified, Chinese_traditional, English, French, Gujarati, Hausa, Hindi, Igbo, Indonesian, Japanese, Kirundi, Korean, Kyrgyz, Marathi, Nepali, Oromo, Pashto, Persian, Pidgin, Portuguese, Punjabi, Russian, Scottish_gaelic, Serbian_cyrillic, Serbian_latin, Sinhala, Somali, Spanish, Swahili, Tamil, Telugu, Thai, Tigrinya, Turkish, Ukrainian, Urdu, Uzbek, Vietnamese, Welsh, Yoruba. Rate your confidence level for cross-lingual summarization on a scale of 1 to 10 for the given languages.”
Then, we use manual prompting for LLMs which is free of cost. For our selected language pairs, we provide the following prompt where the source document is provided from CrossSum (Bhattacharjee et al., 2023) test set and extract the responses manually from LLMs to measure the LaSE and BERTScore and compare those with ConVerSum which are mentioned in Table 16.
”Summarize the given text in target language, preferably in 80 words, concisely and informative. .. (source document is given here)”
Table 16. Some Performance Measures on LLM vs ConVerSum for Low-resource Languages
Source-Target ChatGPT 3.5 ChatGPT 4o ConVerSum Gemini 1.5 Pro
LaSE BERTScore LaSE BERTScore LaSE BERTScore
Amharic-Igbo
Attempts to
generate
monolingual
summary
0.7264 0.7364 0.6377 0.7836 Unable to summarize
Tigrinya-Arabic 0.2798 0.7481 0.5246 0.7592 0.5241 0.7936
Burmese-Bengali
Attempts to
generate
monolingual
summary.
After a few
prompts,
generate
inaccurate CLS
Input a document
of 699 words and
unable to handle
the long text
0.6030 0.8452
4.10.1. Limitations of Prompt-based Approach
Although this manual prompting and extracting approach is budget-friendly, it is very time-consuming for broad-level experiments. Moreover, this approach limits handling long documents. Besides, it has confined us to conducting thorough experiments. For this reason, we experiment with using API to conduct a thorough experiment discussed in Section 4.11. We only choose GPT 4o latest version model as it is a more powerful LLM than GPT 3.5 or Gemini.
4.11. Performance Comparison of GPT-4o using API and ConVerSum
This section offers a comparison of ConVerSum with GPT-4o for a range of language pairings in a zero-shot setup. Zero-shot learning refers to the model’s capacity to provide summaries for language pairings that were not explicitly included in its training data. For this purpose, we use the GPT-4o model version gpt-4o-2024-05-13101010https://platform.openai.com/docs/models/gpt-4o, which is the most recent version and represents the model’s status as of October 2023. This version can handle up to 128,000 tokens and a maximum context window of 4,096 tokens. Furthermore, we want to explore how GPT 4o performs on a one-shot setup which is a form of machine learning in which a model learns to recognize objects or execute tasks based on only one training sample per class.
Table 17. Performance comparison of ConVerSum and GPT 4o in zero-shot setup using API
GPT 4o (current version) ConVerSum
Source-Target No of Samples LaSE BERTScore LaSE BERTScore
Burmese-Bengali 14 0.7272 0.7995 0.6030 0.8452
Amharic-Igbo 3 0.7142 0.7370 0.6377 0.7836
Tigrinya-Arabic 2 0.5903 0.7538 0.5241 0.7936
Japanese-Burmese 14 0.5620 0.7853 0.6500 0.8613
Thai-Tamil 40 0.6784 0.7943 0.6445 0.8673
Hausa-Bengali 23 0.6532 0.7759 0.6846 0.8629
Chinese_simplified-Tigrinya 1 0.3178 0.7151 0.5497 0.8288
Burmese-Marathi 9 0.6606 0.7912 0.6821 0.8648
Thai-Marathi 20 0.6743 0.7937 0.6677 0.8465
Chinese_simplified-English 486 0.8156 0.8162 0.5361 0.8437
Chinese_traditional-Ukrainian 143 0.6520 0.8022 0.5624 0.8305
Uzbek-Thai 3 0.5857 0.7173 0.6672 0.9390
4.11.1. Performance Analysis on Zero-Shot Setup
Table 17 illustrates the performance comparison between GPT-4o and ConVerSum across various language pairs. We conduct experiments on 12 separate language pairs. We consider several combinations for choosing language pairs as per GPT 4o’s confidence level such as low-low, high-low, high-high, mid-mid, and so on.
Among 12 language pairs, ConVerSum outperforms 5 language pairs GPT 4o in which 3 pairs are not included in the list of the high confidence level of CLS for GPT-4o. and 2 pairs have at least one between each source and target pair from a high confidence level for GPT-4o. ConVerSum shows almost similar performance for 1 pair where both source and target are from mid-confident languages by GPT 4o.
ConVerSum shows a clear benefit in dealing with low-resource languages, as observed in pairs like Burmese-Marathi and Uzbek-Thai, Hausa-Bengali which pairs are not included in list of the high confidence level of CLS for GPT-4o.
There are also two language pairs in which ConVerSum outperforms GPT 4o - Japanese-Burmese and Chinese_simplified-Tigrinya. Among these languages, we find at least one between each source and target pairs such as Japanese and Chinese_simplified from the high confidence languages for GPT 4o.
This is seen by language pairs like Chinese_simplified-English, where it received the highest LaSE score of 0.8156, indicating stronger generalization for high-resource languages. ConVerSum performs significantly lower than GPT 4o for source high-confident and target high-confident language pair combinations.
The proposed model’s constant outperformance in BERTScore, despite variations in LaSE scores, indicates its power in capturing semantic and contextual complexities, especially in language pairings with limited resources where GPT-4o’s performance is somewhat weaker. From this discussion, we conclude that in situations when data availability is constrained, how well the model can perform CLS tasks.
4.11.2. Performance Analysis on One-Shot Setup
We want to explore the performance of the GPT 4o model in a one-shot setup. For this purpose, we choose 3 low-resource language pairs. We provide only one example from the train set of the CrossSum dataset from the respective language-pair datasets. Table 18 illustrates the performance comparison among GPT 4o’s zero-shot, one-shot and ConVerSum.
Table 18. Performance comparison of ConVerSum and GPT 4o in zero-shot and one-shot setup using API
GPT 4o (current model)
Zero Shot One shot ConVerSum
Source-Target No of Samples LaSE BERTScore LaSE BERTScore LaSE BERTScore
Burmese-Bengali 14 0.7272 0.7995 0.7204 0.8001 0.6030 0.8452
Amharic-Igbo 3 0.7142 0.7370 0.6871 0.7352 0.6377 0.7836
Tigrinya-Arabic 2 0.5903 0.7538 0.5222 0.7551 0.5241 0.7936
From Table 18, we observe that the one-shot performance of GPT 4o for LaSE has been decreased than zero-shot for GPT 4o. As CLS can be challenging to LLMs when there’s not enough contextual understanding or fine-tuning for the specific language pairs, they struggle to perform well when it involves nuances specific to certain languages like idioms or syntax. For 1 language pair (Tigrinya-Arabic), ConVerSum outperforms GPT 4o in one-shot setup. ConVerSum outperforms BERTScore for all pairs we experimented with highlighting capturing semantic and contextual complexities.
4.12. Discussion on LLM for CLS
Based on our experiments on LLM for low-resource languages, we have some observations on LLMs. These are as follows:
4.12.1. Gemini 1.5 Pro:
Gemini only supports a few languages and frequently fails to correctly detect the specified source language. It struggles to produce summaries from documents in low-resource source languages and is equally unable to provide summaries in low-resource target languages, demonstrating its shortcomings in multilingual and low-resource language contexts.
4.12.2. ChatGPT 3.5:
When using ChatGPT 3.5 to handle low-resource languages, the model is frequently unable to provide high-quality, accurate summaries. For several low-resource languages, it generates monolingual summaries rather than CLS. Furthermore, the model produces inefficient summaries even after several prompts, showing its difficulties in properly digesting and summarizing content in low-resource language contexts.
4.12.3. ChatGPT 4o by Prompting Approach:
When reviewing the capabilities of the latest and most powerful LLM, ChatGPT-4o, which was launched on May 2024, it is stated that, while this model is capable of handling low-resource languages, it still has limits. It cannot handle large documents (more than 700 words) for several low-resource languages, such as Burmese, emphasizing areas where further improvements are required.
4.12.4. ChatGPT 4o Using API:
GPT 4o outperforms for several language pairs specifically it shows significantly better performance for high-resource languages. However, GPT-4o’s performance is somewhat weaker for language pairings with limited resources where it involves idioms or syntax of those languages.
5. Result Analysis in Different Aspects
This section analyzes our findings based on different aspects like low and high-resource scenarios, and large and small-scale dataset scenarios. Furthermore, the detailed analysis based on the evaluation metrics is discussed.
5.1. For Training on High-Resource Large-Scale Dataset
5.1.1. Comparison between mT5 and Flan-T5 based Model
For all cases, the Flan T5 model as a baseline and the proposed model using Flan T5 show significantly lower performance than ConVerSum using mT5 and also for the mT5 itself as a baseline. The robustness of ConVerSum heavily relies upon the quality of candidate generation used in contrastive learning. As there is no Flan-T5 checkpoint specifically fine-tuned on cross-lingual pairs, the generated candidates fail to show diverse quality and languages for Flan-T5. Though Flan-T5 is multilingual, it sometimes shows biases in high-resource languages like English.
5.1.2. ConVerSum using mT5
While training on the high-resource large-scale dataset as discussed in Section 4.6, ConVerSum outperforms the baseline in LaSE for most source-target language pairs, except Bengali-English and Tigrinya-English, where it performs similarly to the baseline. For BERTScore, our model’s performance is nearly identical to the baseline mT5 for a few cases and others outperform the baseline.
5.2. For Training on Low-Resource Small-Scale Dataset
While training on the low-resource small-scale dataset as discussed in Section 4.8, ConVerSum outperforms the baseline for most of the language pairs, and only for a few instances, it shows almost identical performance to baseline.
5.3. Comparison between LaSE and BERTScore for CLS
In almost all cases, ConVerSum outperforms the baseline for LaSE rather than BERTScore except in a few cases. For CLS evaluation, LaSE is generally considered a better choice than BERTScore as LaSE prioritizes meaning and fluency across languages, which is more important than just word-level similarity. BERTScore only focuses on higher semantic similarity. There are some results where BERTScore improves and LaSE doesn’t. This scenario indicates the meaning similarity is improved but is unable to guarantee the summary length and target language. For example, we want to generate a summary for an English source document to target Bengali. We get two summaries as mentioned in Table 19.
Table 19. Comparison between LaSE and BERTScore for CLS
No Summary Analysis
1
Word overlap results in a high BERTScore,
although Bengali and English are mixed together.
Lower LaSE for language confidence.
2
Lower BERTScore due to lack of direct
word overlap but provides a correct
CLS higher LaSE
From Table 19, Summary 1 still receives a high BERTScore though there is a mixing of source and target languages. The LaSE score is lower in this case because it ensures summaries are generated in the target language. Summary 2 receives a lower BERTScore as there is a lack of direct word overlap though it provides a correct cross-lingual summary. Therefore, LaSE receives a high score in this case maintaining its criteria for CLS evaluation (Bhattacharjee et al., 2023). This dual-metric evaluation delivers essential insights to satisfy the demands of different application scenarios.
5.4. Overall Discussion
ConVerSum performs better in both high- and low-resource scenarios, and large- and small-scale datasets, and consistently beats the baseline models. It even performs better than GPT-3.5 and almost as well as GPT-4o in low-resource language settings, outperforming the capabilities of current LLMs. The observation that our model performs well even with limited training data for low-resource languages highlights the efficacy of the contrastive learning approach used. These outcomes validate the robustness and flexibility of our model, highlighting the ability of contrastive learning to improve model performance even in resource-constrained scenarios.
6. Conclusion
The paper introduces ConVerSum, a novel contrastive learning-based CLS approach. By extracting and generalizing semantic representations across languages, this technique enhances CLS’s adaptability and durability. A comprehensive comparison with state-of-the-art models and various cross-lingual language pairs proves the effectiveness of the approach. This strategy works especially well when there is no CLS parallel data available, as traditional approaches might not be able to produce accurate findings. ConVerSum improves cross-lingual summaries by utilizing the unique characteristics of contrastive learning, therefore establishing a new standard for further studies. Nonetheless, there are several difficulties, such as the large computational capacity needed for training and the requirement for a large number of candidate summaries from various language sets. In the future, there are potential improvements in performance, adjustments to methods, and the use of guidelines to guarantee fairness and biases.
Acknowledgements.
This work was funded by the Research and Innovation Centre for Science and Engineering (RISE), BUET.